OpenStudy (anonymous):

system of equations 8(x)+10(y)=152.50 12(x)+8(y)=171.00 what is x or y

OpenStudy (anonymous):

this is what i have so far X= pet store Y= movie 8(x)+10(y)=152.50 12(x)+8(y)=171.00 12(x)+8(y)=171.00- 8(x)+10(y)=152.50 4x-2y= 18.50 2x-y=9.25 2x-9.25=y and now im lost

OpenStudy (anonymous):

\[\left[\begin{matrix}8 & 10 & 152.50\\ \\ 12& 8 &171\end{matrix}\right] \] using RREF ull get \[\left[\begin{matrix}1 & 0 &8.75 \\ 0 & 1 &8.25\end{matrix}\right]\]

OpenStudy (anonymous):

therefore x=8.75 and y =8.25

OpenStudy (anonymous):

im lost that doesn't make any sense to me

OpenStudy (anonymous):

wait give a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ive never heard of this method how did you get the 8 10 12 8 to change to 1 0 1 0?

OpenStudy (anonymous):

its the reduced row echelon form

OpenStudy (anonymous):

but you can also solve it using elimination method

OpenStudy (anonymous):

using elimination process multiply the first equation by -8 and the second by 10